OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <10, 0>, v = <0, -9>

OpenStudy (solomonzelman):

If, \(\color{#000000 }{ \displaystyle {\bf u}\cdot {\bf v}=0 }\) Then, your two vectors are perpendicular (orthogonal). If, \(\color{#000000 }{ \displaystyle {\bf u}\times {\bf v}=0 }\) Then, your two vectors are parallel. And if niether of these conditions are met, then they are not perpendicular and not parallel (i.e. niether).

OpenStudy (anonymous):

Oh alright. That helps. Difference between u⋅v=0 and u×v=0 mainly ⋅ vs x i think im forgetting what ⋅ represents

OpenStudy (solomonzelman):

• is a dot product × is a vector product, or the cross product

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle {\bf a}=\left(a_x,~a_y\right) }\) \(\color{#000000 }{ \displaystyle {\bf b}=\left(b_x,~b_y\right) }\) \(\color{#000000 }{ \displaystyle {\bf a}\cdot {\bf b}=a_x{\tiny~}b_x+ a_y{\tiny~}b_y}\)

OpenStudy (solomonzelman):

does this post make sense, or is it something rediculous?

OpenStudy (anonymous):

Ahh okay. that makes sense. Thank you solomon

OpenStudy (solomonzelman):

Ok, when you have the answer, you can post it here if you feel like, and I will verify it.

OpenStudy (solomonzelman):

Good Luck!

OpenStudy (anonymous):

I'm thinking its orthogonal. Thank you solomon for the help.

OpenStudy (solomonzelman):

yes, they are perpendicular, good job!